# Positivity is undecidable in tensor products of free algebras

### 09 Dec 2023 v1

@@ 2.3. Group theory
In addition to the elementary definitions above, we use some more advanced concepts from group theory. Suppose $S$ is a finite set of generators for a group $G$, and let $\mathcal{L}$ be the set of words over $S\cup S^{-1}$ which are trivial in $G$. The word problem for $G$ is the problem of determining whether a word over $S\cup S^{-1}$ belongs to $\mathcal{L}$. Accordingly, we say that the word problem for $G$ with respect to $S$ is solvable in non-deterministic time $T:\mathbb{N}\rightarrow\mathbb{N}$ (resp. is in NP) if $\mathcal{L}$ can be recognized in time $T$ (resp. in polynomial time).
@@
Lemma 2.2 (Britton’s lemma, [Rot73]).
@@
Theorem 2.4 ([BORS02]).
Let $G$ be a group with finite generating set $S$. If the word problem for $G$ with respective to $S$ is solvable in non-deterministic time $T(n)$, where $T^{4}$ is superadditive, then $G$ can be embedded into a finitely presented group $H$ with isoperimetric function $n^{2}T(n^{2})^{4}$.
Here a function $T:\mathbb{N}\rightarrow\mathbb{N}$ is superadditive if $T(m+n)\geq T(m)+T(n)$ for all $m,n$. It follows immediately from this theorem that if the word problem for $G$ is in NP, then $G$ can be embedded in a finitely presented group $H$ with a polynomial isoperimetric function.
@@ 3. Approximate representations
Let $\mathcal{A}$ be a finitely generated ***-algebra. Suppose $\epsilon\geq 0$ and $\mathcal{R}\subseteq\mathcal{A}$. An $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$ is a state $\varphi$ on $\mathcal{A}$ such that $\varphi(r^{*}r)\leq\epsilon$ for all $r\in\mathcal{R}\cup\mathcal{R}^{*}$. If, in addition, $\varphi$ is tracial, then we say that $\varphi$ is a tracial $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$.
If $(\pi:\mathcal{A}\to\mathcal{B}(\mathcal{H}),\xi)$ is the GNS representation of a state $\varphi$ on $\mathcal{A}$, then $\lVert x\rVert_{\xi}:=\lVert x\cdot\xi\rVert$ is a semi-norm on $\mathcal{B}(\mathcal{H})$ such that $\lVert a\rVert_{\varphi}=\lVert\pi(a)\rVert_{\xi}$ for all $a\in\mathcal{A}$. If $\varphi$ is an $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$ as in Definition 3.1, then $\lVert\pi(r)\rVert_{\xi}\leq\sqrt{\epsilon}$ for all $r\in\mathcal{R}$, so we can think of $\pi$ as a $\sqrt{\epsilon}$-approximate representation for *$\mathbb{C}^{*}\langle\mathcal{A}:\mathcal{R}\rangle$ with respect to $\lVert\cdot\rVert_{\xi}$.
@@
[…]
@@
    $u_{i},v_{i}$ are ***-monomials in *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ for all $1\leq i\leq n$,
@@
Here a ***-monomial in *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is a product $a_{1}a_{2}\cdots a_{k}$ where $k\geq 0$ and $a_{1},\ldots,a_{k}\in\mathcal{X}\cup\mathcal{X}^{*}$ (the integer $k$ is called the degree). We let $\lVert\cdot\rVert_{1}$ and $\lVert\cdot\rVert_{1,1}$ denote the $1$-norm and first Sobolev $1$-seminorm on *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ with respect to the basis of ***-monomials. In other words, if $f=\sum_{i=1}^{n}\lambda_{i}u_{i}\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ for distinct ***-monomials $u_{1},\ldots,u_{n}$, then $\lVert f\rVert_{1}:=\sum_{i=1}^{n}|\lambda_{i}|$ and $\lVert f\rVert_{1,1}:=\sum_{i=1}^{n}|\lambda_{i}|\deg(u_{i})$. Note that $\lVert\cdot\rVert_{1}$ is submultiplicative, and that $\lVert f^{*}\rVert_{1}=\lVert f\rVert_{1}$, $\lVert f^{*}\rVert_{1,1}=\lVert f\rVert_{1,1}$ for all $f\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$. The following elementary lemma is useful for computing sizes of $\mathcal{R}$-decompositions.
@@
Let $\mathcal{A}$ be a ***-algebra with generating set $\mathcal{X}$, and let *$\mathcal{R}\subseteq\mathbb{C}^{*}\langle\mathcal{X}\rangle$ be a set of ***-polynomials over $\mathcal{X}$. Then:
@@
    If $f=\sum_{i=1}^{n}u_{i}r_{i}v_{i}$ for ***-polynomials $u_{1},\ldots,u_{n},v_{1}\ldots,v_{n}\in\mathbb{C}^{*}\langle X\rangle$ and $r_{1},\ldots,r_{n}\in\mathcal{R}\cup\mathcal{R}^{*}$, then $f$ has an $\mathcal{R}$-decomposition of size at most
@@
    If $f-g$ has an $\mathcal{R}$-decomposition of size $\leq\Lambda_{1}$, and $g-h$ has an $\mathcal{R}$-decomposition of size $\leq\Lambda_{2}$, then $f-h$ has an $\mathcal{R}$-decomposition of size $\leq\Lambda_{1}+\Lambda_{2}$.
@@
[…]
If $\mathcal{X}$ is a set of unitary generators for $\mathcal{A}$, then all ***-monomials in *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ are unitary in $\mathcal{A}$. As a result, an $\mathcal{R}$-decomposition of $f$ can be used to bound the state-dependent norm of $f$ with respect to any tracial $(\epsilon,\mathcal{R})$-state:
@@
[…]
@@
The proof of Proposition 3.4 does not work if $\tau$ is not tracial. However, we can still bound $\lVert f\rVert_{\tau}$ using the size of an $\mathcal{R}$-decomposition if $\tau$ is the restriction of a state on $\mathcal{A}\otimes\mathcal{A}$ satisfying a synchronicity condition:
@@
Suppose $\mathcal{A}$ is a ***-algebra. A state $\varphi$ on $\mathcal{A}\otimes\mathcal{A}$ is $(\epsilon,\mathcal{X})$-synchronous for some $\epsilon\geq 0$ and $\mathcal{X}\subseteq\mathcal{A}$ if
@@
When working with $(\epsilon,\mathcal{X})$-synchronous states, it is convenient to let $\omega_{\mathcal{X}}$ denote the linear map *$\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathbb{C}^{*}\langle\mathcal{X}\rangle$ sending $x_{1}x_{2}\cdots x_{n}\mapsto x_{n}x_{n-1}\cdots x_{1}$ for all $n\in\mathbb{N}$ and $x_{1},\ldots,x_{n}\in\mathcal{X}\cup\mathcal{X}^{*}$. This map satisfies $\omega_{\mathcal{X}}(a^{*})=\omega_{\mathcal{X}}(a)^{*}$ and $\omega_{\mathcal{X}}(ab)=\omega_{\mathcal{X}}(b)\omega_{\mathcal{X}}(a)$ for all $a,b\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$, and hence $\omega_{\mathcal{X}}$ is a ***-isomomorphism *$\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathbb{C}^{*}\langle\mathcal{X}\rangle^{\mathrm{op}}$, where *$\mathbb{C}^{*}\langle\mathcal{X}\rangle^{\mathrm{op}}$ is the opposite algebra of *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$.
@@
Let $\mathcal{A}$ be a ***-algebra generated by a finite set of unitaries $\mathcal{X}$. If $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$, then
@@
Thus Equation 3.1 holds for all $x\in\mathcal{X}\cup\mathcal{X}^{*}$. Suppose $u=x_{1}\cdots x_{n}$ is a monomial in *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ of degree $n\geq 1$. Let $x_{0}=x_{n+1}=1$. Then
@@
Let $\mathcal{A}$ be a ***-algebra generated by a finite set of unitaries $\mathcal{X}$, and let $\varphi$ be an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$. If $u$ is a monomial in *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $a\in\mathcal{A}$, then $|\lVert u^{*}au\otimes 1\rVert_{\varphi}-\lVert a\otimes 1\rVert_{\varphi}|\leq\lVert a\rVert_{\mathcal{A}}\deg(u)\sqrt{\epsilon}$.
@@
Since $u$ is a monomial in *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$, $u$ and $\omega_{\mathcal{X}}(u)$ are unitary in $\mathcal{A}$. By Lemma 3.6, $\lVert u\otimes 1-1\otimes\omega_{\mathcal{X}}(u)\rVert_{\varphi}\leq\deg(u)\sqrt{\epsilon}$. Thus
@@
Let $\mathcal{A}$ be a ***-algebra generated by a finite set of unitaries $\mathcal{X}$, and let $\varphi$ be an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$. If $u$ is a monomial in *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $a\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is a hermitian square in $\mathcal{A}$, then
@@
Let $\mathcal{A}$ be a ***-algebra generated by a finite set of unitaries $\mathcal{X}$, and let $\iota:\mathcal{A}\to\mathcal{A}\otimes\mathcal{A}:a\mapsto a\otimes 1$ be the left inclusion. Suppose *$\mathcal{R}\subseteq\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is a set of ***-polynomials over $\mathcal{X}$, and let $\varphi$ be an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$ such that $\tau:=\varphi\circ\iota$ is an $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$. If $f\in\mathcal{A}$ has an $\mathcal{R}$-decomposition of size $\Lambda$, then $\lVert f\rVert_{\tau}\leq\Lambda\sqrt{\epsilon}$.
@@
Let $\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}$ be an $\mathcal{R}$-decomposition for $f$ in $\mathcal{A}$ of size $\Lambda$. For every $1\leq i\leq n$, by Lemma 3.6 there is a monomial $\widehat{v}_{i}$ in *$\mathbb{C}^{*}\langle\mathcal{X}\rangle$ such that $\lVert v_{i}\otimes 1-1\otimes\widehat{v}_{i}\rVert_{\varphi}\leq deg(v_{i})\sqrt{\epsilon}$. Since $\varphi\circ\iota$ is an $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$ and $1\otimes\widehat{v}$ is unitary, it follows that $\lVert r_{i}\otimes\widehat{v}_{i}\rVert_{\varphi}=\lVert r_{i}\rVert_{\varphi\circ\iota}\leq\sqrt{\epsilon}$ for all $1\leq i\leq n$. Then
@@
$\leq\sum_{i=1}^{n}\lvert\lambda_{i}\rvert\big{(}\lVert(r_{i}\otimes 1)(v_{i}\otimes 1-1\otimes\widehat{v}_{i})\rVert_{\varphi}+\lVert r_{i}\otimes\widehat{v}_{i}\rVert_{\varphi}\big{)}$

### 18 Mar 2026 v2

@@ 2.3. Group theory
In addition to the elementary definitions above, we use some more advanced concepts from group theory. Suppose $S$ is a finite set of generators for a group $G$, and let $\mathcal{L}$ be the set of words over $S\cup S^{-1}$ which are trivial in $G$. The word problem for $G$ is the problem of determining whether a word over $S\cup S^{-1}$ belongs to $\mathcal{L}$. Accordingly, we say that the word problem for $G$ with respect to $S$ is solvable in non-deterministic time $T:\mathbb{N}\rightarrow\mathbb{N}$ (resp. non-deterministic polynomial time) if $\mathcal{L}$ can be recognized in time $T$ (resp. polynomial time).
@@
Lemma 2.2 (Britton’s lemma, Theorem 11.81 of [ROT73]).
@@
Theorem 2.4 ([BOR+02]).
Let $G$ be a group with finite generating set $S$. If the word problem for $G$ with respect to $S$ is solvable in non-deterministic time $T(n)$, where $T^{4}$ is superadditive, then $G$ can be embedded into a finitely presented group $H$ with isoperimetric function $n^{2}T(n^{2})^{4}$.
Here a function $T:\mathbb{N}\rightarrow\mathbb{N}$ is superadditive if $T(m+n)\geq T(m)+T(n)$ for all $m,n$. It follows immediately from this theorem that if the word problem for $G$ can be solved in non-deterministic polynomial time, then $G$ can be embedded in a finitely presented group $H$ with a polynomial isoperimetric function.
@@ 3. Approximate representations
Let $\mathcal{A}$ be a finitely generated *-algebra. Suppose $\epsilon\geq 0$ and $\mathcal{R}\subseteq\mathcal{A}$. An $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$ is a state $\varphi$ on $\mathcal{A}$ such that $\varphi(r^{*}r)\leq\epsilon$ for all $r\in\mathcal{R}\cup\mathcal{R}^{*}$. If, in addition, $\varphi$ is tracial, then we say that $\varphi$ is a tracial $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$.
If $(\pi:\mathcal{A}\to\mathcal{B}(\mathcal{H}),\xi)$ is the GNS representation of a state $\varphi$ on $\mathcal{A}$, then $\lVert x\rVert_{\xi}:=\lVert x\cdot\xi\rVert$ is a semi-norm on $\mathcal{B}(\mathcal{H})$ such that $\lVert a\rVert_{\varphi}=\lVert\pi(a)\rVert_{\xi}$ for all $a\in\mathcal{A}$. If $\varphi$ is an $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$ as in Definition 3.1, then $\lVert\pi(r)\rVert_{\xi}\leq\sqrt{\epsilon}$ for all $r\in\mathcal{R}$, so we can think of $\pi$ as a $\sqrt{\epsilon}$-approximate representation for $\mathbb{C}^{*}\langle\mathcal{A}:\mathcal{R}\rangle$ with respect to $\lVert\cdot\rVert_{\xi}$.
@@
Let $\mathcal{A}$ be a *-algebra with generating set $\mathcal{X}$. Let $\mathcal{R}\subseteq\mathbb{C}^{*}\langle\mathcal{X}\rangle$ be a set of *-polynomials over $\mathcal{X}$. For any *-polynomial $f\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ that is trivial in $\mathcal{A}/\langle\mathcal{R}\rangle$, we say that $\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}$ is an $\mathcal{R}$-decomposition for $f$ in $\mathcal{A}$ if
@@
    $u_{i},v_{i}$ are *-monomials in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ for all $1\leq i\leq n$,
@@
Here $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ means the set of formal *-polynomials with indeterminates $\mathcal{X}$. There is a natural map from $\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathcal{A}$ sending $x\in\mathcal{X}$ to its image in $\mathcal{A}$, so we can also regard elements of $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ as elements of $\mathcal{A}$, e.g. when taking the quotient $\mathcal{A}/\langle\mathcal{R}\rangle$. A *-monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is a product $a_{1}a_{2}\cdots a_{k}$ where $k\geq 0$ and $a_{1},\ldots,a_{k}\in\mathcal{X}\cup\mathcal{X}^{*}$ (the integer $k$ is called the degree).
We let $\lVert\cdot\rVert_{1}$ and $\lVert\cdot\rVert_{1,1}$ denote the $1$-norm and first Sobolev $1$-seminorm on $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ with respect to the basis of *-monomials. In other words, if $f=\sum_{i=1}^{n}\lambda_{i}u_{i}\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ for distinct *-monomials $u_{1},\ldots,u_{n}$, then $\lVert f\rVert_{1}:=\sum_{i=1}^{n}|\lambda_{i}|$ and $\lVert f\rVert_{1,1}:=\sum_{i=1}^{n}|\lambda_{i}|\deg(u_{i})$. Note that $\lVert\cdot\rVert_{1}$ is submultiplicative, and that $\lVert f^{*}\rVert_{1}=\lVert f\rVert_{1}$, $\lVert f^{*}\rVert_{1,1}=\lVert f\rVert_{1,1}$ for all $f\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$. The following elementary lemma is useful for computing sizes of $\mathcal{R}$-decompositions.
@@
Let $\mathcal{A}$ be a *-algebra with generating set $\mathcal{X}$, and let $\mathcal{R}\subseteq\mathbb{C}^{*}\langle\mathcal{X}\rangle$ be a set of *-polynomials over $\mathcal{X}$. Then:
@@
    If $f=\sum_{i=1}^{n}u_{i}r_{i}v_{i}$ for *-polynomials $u_{1},\ldots,u_{n},v_{1}\ldots,v_{n}\in\mathbb{C}^{*}\langle X\rangle$ and $r_{1},\ldots,r_{n}\in\mathcal{R}\cup\mathcal{R}^{*}$, then $f$ has an $\mathcal{R}$-decomposition of size at most
@@
    If fgf-g has an $\mathcal{R}$-decomposition of size $\leq\Lambda_{1}$, and ghg-h has an $\mathcal{R}$-decomposition of size $\leq\Lambda_{2}$, then fhf-h has an $\mathcal{R}$-decomposition of size $\leq\Lambda_{1}+\Lambda_{2}$.
@@
For instance, suppose we can demonstrate that $g=h$ in $\mathcal{A}/\langle\mathcal{R}\rangle$ by finding $f_{1},\ldots,f_{n}$, $u_{1},\ldots,u_{n}$, $v_{1},\ldots,v_{n}\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $r_{1},\ldots,r_{n}\in\mathcal{R}\cup\mathcal{R}^{*}$ such that $f_{i+1}-f_{i}=u_{i}r_{i}v_{i}$ for all $0\leq i\leq n$, where $f_{0}=g$ and $f_{n+1}=h$. Then we know that ghg-h has an $\mathcal{R}$-decomposition of size at most $\sum_{i=1}^{n}\lVert u_{i}\rVert_{1}\lVert v_{i}\rVert_{1}+\lVert u_{i}\rVert_{1}\lVert r_{i}\rVert_{\mathcal{A}}\lVert v_{i}\rVert_{1,1}$.
If $\mathcal{X}$ is a set of unitary generators for $\mathcal{A}$, then all *-monomials in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ are unitary in $\mathcal{A}$. As a result, an $\mathcal{R}$-decomposition of $f$ can be used to bound the state-dependent norm of $f$ with respect to any tracial $(\epsilon,\mathcal{R})$-state:
@@
Suppose $\mathcal{A}$ is a *-algebra generated by a finite set of unitaries $\mathcal{X}$. Let $\mathcal{R}\subset\mathbb{C}^{*}\langle\mathcal{X}\rangle$ be a set of *-polynomials over $\mathcal{X}$, and let $\tau$ be a tracial $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$. If $f\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is trivial in $\mathcal{A}/\langle\mathcal{R}\rangle$ with an $\mathcal{R}$-decomposition $\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}$, then $\lVert f\rVert_{\tau}\leq(\sum_{i=1}^{n}|\lambda_{i}|)\sqrt{\epsilon}$.
@@
The proof of Proposition 3.4 does not work if $\tau$ is not tracial. However, we can still bound $\lVert f\rVert_{\tau}$ using the size of an $\mathcal{R}$-decomposition if $\tau$ is the restriction of a state on $\mathcal{A}\otimes\mathcal{A}$ satisfying a synchronicity [^2] The term synchronous is borrowed from nonlocal games in quantum information. Roughly speaking, a state $\varphi$ is synchronous if the action of $x\in\mathcal{X}$ on $\varphi$ is the same in both tensor factors. condition:
@@
Suppose $\mathcal{A}$ is a *-algebra. A state $\varphi$ on $\mathcal{A}\otimes\mathcal{A}$ is $(\epsilon,\mathcal{X})$-synchronous for some $\epsilon\geq 0$ and $\mathcal{X}\subseteq\mathcal{A}$ if
@@
When working with $(\epsilon,\mathcal{X})$-synchronous states, it is convenient to let $\omega_{\mathcal{X}}$ denote the linear map $\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathbb{C}^{*}\langle\mathcal{X}\rangle$ sending $x_{1}x_{2}\cdots x_{n}\mapsto x_{n}x_{n-1}\cdots x_{1}$ for all $n\in\mathbb{N}$ and $x_{1},\ldots,x_{n}\in\mathcal{X}\cup\mathcal{X}^{*}$. This map satisfies $\omega_{\mathcal{X}}(a^{*})=\omega_{\mathcal{X}}(a)^{*}$ and $\omega_{\mathcal{X}}(ab)=\omega_{\mathcal{X}}(b)\omega_{\mathcal{X}}(a)$ for all $a,b\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$, and hence $\omega_{\mathcal{X}}$ is a *-isomomorphism $\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathbb{C}^{*}\langle\mathcal{X}\rangle^{\mathrm{op}}$, where $\mathbb{C}^{*}\langle\mathcal{X}\rangle^{\mathrm{op}}$ is the opposite algebra of $\mathbb{C}^{*}\langle\mathcal{X}\rangle$.
@@
Let $\mathcal{A}$ be a *-algebra generated by a finite set of unitaries $\mathcal{X}$. If $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$, then
@@
Thus Equation 3.1 holds for all $x\in\mathcal{X}\cup\mathcal{X}^{*}$. Suppose $u=x_{1}\cdots x_{n}$ is a monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ of degree $n\geq 1$. Let $x_{0}=x_{n+1}=1$. Then
@@
Let $\mathcal{A}$ be a *-algebra generated by a finite set of unitaries $\mathcal{X}$, and let $\varphi$ be an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$. If $u$ is a monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $a\in\mathcal{A}$, then $|\lVert u^{*}au\otimes 1\rVert_{\varphi}-\lVert a\otimes 1\rVert_{\varphi}|\leq\lVert a\rVert_{\mathcal{A}}\deg(u)\sqrt{\epsilon}$.
@@
Since $u$ is a monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$, $u$ and $\omega_{\mathcal{X}}(u)$ are unitary in $\mathcal{A}$. By Lemma 3.6, $\lVert u\otimes 1-1\otimes\omega_{\mathcal{X}}(u)\rVert_{\varphi}\leq\deg(u)\sqrt{\epsilon}$. Thus
@@
Let $\mathcal{A}$ be a *-algebra generated by a finite set of unitaries $\mathcal{X}$, and let $\varphi$ be an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$. If $u$ is a monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $a\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is a hermitian square in $\mathcal{A}$, then
@@
Let $\mathcal{A}$ be a *-algebra generated by a finite set of unitaries $\mathcal{X}$, and let $\iota:\mathcal{A}\to\mathcal{A}\otimes\mathcal{A}:a\mapsto a\otimes 1$ be the left inclusion. Suppose $\mathcal{R}\subseteq\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is a set of *-polynomials over $\mathcal{X}$, and let $\varphi$ be an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$ such that $\tau:=\varphi\circ\iota$ is an $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$. If $f\in\mathcal{A}$ has an $\mathcal{R}$-decomposition of size $\Lambda$, then $\lVert f\rVert_{\tau}\leq\Lambda\sqrt{\epsilon}$.
@@
Let $\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}$ be an $\mathcal{R}$-decomposition for $f$ in $\mathcal{A}$ of size $\Lambda$. By Lemma 3.6, $\lVert v_{i}\otimes 1-1\otimes\omega_{\mathcal{X}}(v_{i})\rVert_{\varphi}\leq deg(v_{i})\sqrt{\epsilon}$ for every $1\leq i\leq n$. Since $\varphi\circ\iota$ is an $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$ and $1\otimes\omega_{\mathcal{X}}(v_{i})$ is unitary in $\mathcal{A}\otimes\mathcal{A}$, it follows that $\lVert r_{i}\otimes\omega_{\mathcal{X}}(v_{i})\rVert_{\varphi}=\lVert r_{i}\rVert_{\varphi\circ\iota}\leq\sqrt{\epsilon}$ for all $1\leq i\leq n$. Then
@@
$\leq\sum_{i=1}^{n}\lvert\lambda_{i}\rvert\big(\lVert(r_{i}\otimes 1)(v_{i}\otimes 1-1\otimes\omega_{\mathcal{X}}(v_{i}))\rVert_{\varphi}+\lVert r_{i}\otimes\omega_{\mathcal{X}}(v_{i})\rVert_{\varphi}\big)$